# Large scale simulations of photosynthetic antenna systems: interplay of cooperativity and disorder

### v1

@@ 2.1 Antenna complexes in Green Sulfur Bacteria (GSB)
Complex A - single cylinder. In Fig.(1 A) we present the model utilized for a single cylinder in the bchQRU triple mutant chlorosome. The cylinder comprises a stack of tightly-packed rings of Bchl c molecules. Each molecule is treated as a dipole with a squared dipole moment , possessing a well defined orientation and position in space 29. Each ring has a radius and contains Bchl molecules separated by a nearest neighbour distance . The rings are equally separated by a vertical distance 29, 51, 38. In our simulations the single cylinder complex can contain up to 10 800 Bchl c molecules, resulting in a maximum length , corresponding to a stack of rings. Further details about the geometry of the bchQRU triple mutant concerning the positions and orientations can be found in Ref. 29.
@@
Complex C - three adjacent concentric cylinders. In panel C of Fig. (1) we show a more elaborate model formed by three adjacent structures, each comprising four concentric cylinders. The wall-to-wall distance between adjacent cylinders is set to nm, a realistic value according to Ref. 24. The maximum length considered here is nm, corresponding to vertical layers in each cylindrical structure. The largest system considered in our simulations contains 132 840 Bchl c molecules, comparable with natural sizes. The positions and orientations of dipoles on each tubular wall are the same as in Complex A and B.
@@
Complex A - light-harvesting system subunit. The RC is directly surrounded by the so-called light harvesting complex I (LHI). LHI is in turn surrounded by several smaller light harvesting complexes, the LHII complexes 2, 53, 4. As shown in panel A of Fig. (2) two RCs are surrounded by the light-harvesting complex LHI B875, which is an S-shaped structure with 56 BChl molecules. The LHI aggregate is a J-aggregate with two superradiant states at 875 nm that are polarized in the ring plane and close to the lowest excitonic state. The LHI aggregate is surrounded by ten LHII rings, each featuring the B850 ring, a J-aggregate composed of 18 BChl molecules with two superradiant states at 850 nm, and the B800 ring composed of 9 BChl molecules with the main absorption peak at 800 nm. This hierarchical structure is able to absorb photons at different frequencies and to funnel the collected energy to the RC 54, 34, 4, 17, 16. In our simulation we model the RC as an aggregate of four BChls: two forming the tightly-coupled special pair and two accessory ones 17. Even if other molecules are present in the RC, these four Bchl molecules are the most relevant for the interaction with the electromagnetic field 55.
Complex B: the chromatophore. We model the whole 3D structure of the chromatophore using data from Ref. 34 collected by AFM, cryo-EM, X-ray crystallography and NMR measurements. In panel B of Fig. (2) our model of the chromatophore is shown with different substructures: i) LHII (blue rings); ii) LHI (orange structures); iii) RC, placed at the center of the LHI structures. Chromatophores, see Fig. (2 B), are connected to the cell membrane at their south pole, therefore the southern polar region is left empty (no molecular aggregates are placed there) to allow the contact with the cell membrane 56. The chromatophore spherical vesicles are reproduced by mapping small planar regions onto spherical ones containing the LHI-RCs or an LHII complex with the area-preserving inverse-Mollweide transformations, see Fig. (LABEL:PBprojection) in LABEL:AppA of the Supporting Information, used also in Ref. 34, 57, 58. While in nature the ratio between LHII and LHI-RC depends on illumination and other conditions, here we consider a vesicle model that contains S-shaped LHI-RCs and LHII complexes. The overall Bchl/RC ratio used in our simulation () is within the natural range for these vesicles, which is between and and corresponds to low-light growth conditions 34 where these systems show the largest efficiency. The data for the LHI, LHII and RC are taken from Ref. 17, 4. Even if it would be more realistic to add random rotations of LHII rings around the axis joining the center of the vesicle and the center of the LHII rings, we checked that the effect of such random rotations is small, see Fig. (LABEL:PBrotationLHII) of LABEL:AppE, and thus we did not include such rotations in our simulations.
@@ 2.3 The radiative Hamiltonian and the dipole approximation
Since photosynthetic antennae operate under natural sunlight, which is very dilute, the single-excitation approximation can be used, so that only states containing a single excitation have been considered. Choosing the basis states in the single-excitation manifold, where represents a state in which the molecule is excited while all the others are in the ground state, the systems can be described through a radiative non-Hermitian Hamiltonian which takes into account the interaction between the molecules mediated by the electromagnetic field (EMF) 13, 12, 59, 12, 60. The radiative Hamiltonian reads:
@@ 2.4 Robustness to thermal noise and static disorder
Given a quantum state specified by the density matrix it is possible to define its coherence length in the single-excitation manifold defined by the basis states 61, 62, 29 in the following way:
@@
is also able to produce a large thermal coherence length at room temperature. Note that even if we consider the coherence length at thermal equilibrium, this does not mean that out-of-equilibrium processes are not important in molecular nanotubes. Indeed in Ref. 63 strong evidence of ultra-fast transport in natural structures with transfer times less than 100 fs have been discussed. Nevertheless thermal equilibrium can be considered as a worst case scenario for coherences. For this reason assuming thermal equilibrium can be considered a good starting point to assess the structural robustness of quantum coherence to thermal noise.
@@
Natural complexes are not only affected by thermal noise but also by other sources of disorder due to the fluctuations in the local environment. In order to analyze the robustness to this kind of disorder, we have considered time-independent and space-dependent fluctuations of the excitation energy of the molecules in the aggregates. Specifically, we consider energy fluctuations which are uniformly distributed around the excitation energy of the molecules , between and , where represents the disorder strength. It is known that static disorder induces eigenmodes localization 64 and quenching of superradiance 65.
@@ 3 Numerical results
In this section we show the numerical results obtained for the antenna complexes of GSB and PB.
@@ 3.1 Green Sulfur Bacteria complexes
Here, the complexes A, B, C of GSB described in Sec. 2.1 have been analyzed. In Fig. (3) the spectrum of the three complexes is shown for the maximal length considered, nm, corresponding to cylinders with 180 layers. Thus, for the three complexes we have (complex A), (complex B), (complex C) chlorophyll molecules. Panels (A-C) in Fig. (3) show that the structure of all complexes allow for the emergence of red-shifted superradiant states, close to the lowest excitonic states. Moreover the amount of maximal superradiance () increases with the system size, see Fig. (4). The largest amount of superradiance is present in complex C which is the largest one and the closest to the natural size and structure of GSB antennae. Note that the fact that in larger aggregates superradiance is enhanced is far from being trivial. Indeed, if we would increase the system size by adding the molecules in the same positions but with randomized dipole directions, no superradiant enhancement would be present, as shown in LABEL:AppF-GSB of the Supporting Information.
@@
The origin of the difference between the NHH and the other Hamiltonian models can be explained by the overlapping resonance criterion 36, see panel D in Fig. (3). Indeed, when the decay width becomes comparable with the eigenmodes mean level spacing, the non-Hermitian part of the Hamiltonian cannot be treated perturbatively and the dipole strength Eq. (6) does not describe anymore the decay widths of the eigenmodes.
@@
We now consider in Fig. (6), the three different complexes for a fixed length corresponding to 100 layers ( nm), smaller than the maximal system length considered in Fig. (3,4). Also in this case we used the HH model to compute both figures of merit since for this length scale the HH model is quite accurate for all complexes, see discussion in LABEL:AppC of the Supporting Information. While the maximal dipole strength decreases with disorder Fig. (6 A), it is still much larger than one even for . On the other hand, the thermal coherence length shows a large robustness to disorder, Fig. (6 B) up to . In Fig. (6 C,D) the maximal dipole strength (panel C) and the thermal coherence length (panel D) are shown at and as a function of the number of chlorophyll molecules present in the three different complexes A,B,C. The most important feature of both figures of merit is that they grow as a larger portion of the photosynthetic antenna is considered, showing that the structure of the GSB photosynthetic antenna as a whole is able to support excitonic coherences. Note that in presence of disorder, the maximal dipole strength grows slower with compared with the case of zero disorder (panel C). On the hand, the value of the coherence length is the same for the two values of disorder strength considered, showing its robustness to disorder (panel D), but as increases it shows a tendency to saturate.
@@ 4 Conclusions
In analyzing different antenna complexes, we compared three distinct Hamiltonian models: the Dipole Hamiltonian, applicable in the small volume limit; the Hermitian Hamiltonian, valid when the decay widths are significantly smaller than the mean level spacing; and the full radiative non-Hermitian Hamiltonian model. For Purple bacteria chromatophores, all models yield consistent results due to the system’s small size (e.g., 60 nm) relative to the absorbed wavelength () and the limited number of Bchl molecules present. However, for the largest complexes considered in Green sulfur bacteria, only the non-Hermitian Hamiltonian model proved effective in providing reliable results.
The presence of very bright (superradiant) states in the whole photosynthetic antenna could shed new light in understanding the functionality of such complexes. In particular, the presence of superradiance and subradiance plays an important role for the efficiency of the energy transfer in these aggregates 66, 17. Let us also remark that the system dynamics strongly depends on the choice of the initial state which is mainly determined by the most superradiant states, see LABEL:AppD for a more detailed discussion about the delocalization of superradiant states in PB and GSB antennae.
@@
The cooperative properties, akin to those observed in photosynthetic aggregates, are inspiring numerous proposals for engineering artificial devices for light-harvesting and clean energy production 67, 68, 69, 70, 71, 72, 73, 74. Thus, our study could significantly impact the development of quantum devices for photon detection and light harvesting.

### v2

@@ 2.1 Antenna complexes in Green Sulfur Bacteria (GSB)
Complex A - single cylinder. In Fig.(1 A) we present the model utilized for a single cylinder in the bchQRU triple mutant chlorosome. The cylinder comprises a stack of tightly-packed rings of Bchl c molecules. Each molecule is treated as a dipole with a squared dipole moment , possessing a well defined orientation and position in space 29. Each ring has a radius and contains Bchl molecules separated by a nearest neighbour distance . The rings are equally separated by a vertical distance 29, 51, 38. This choice of parameters for the geometry of the bchQRU triple mutant type is in agreement with Ref. 38, where the same cylindrical structure is obtained by wrapping the planar lattice under the rolling angle of with respect to the vertical axis of the lattice. In our simulations the single cylinder complex can contain up to 10 800 Bchl c molecules, resulting in a maximum length , corresponding to a stack of rings. Further details about the geometry of the bchQRU triple mutant concerning the positions and orientations can be found in Ref. 29.
@@
Complex C - three adjacent concentric cylinders. In panel C of Fig. (1) we show a more elaborate model formed by three adjacent structures, each comprising four concentric cylinders. The wall-to-wall distance between adjacent cylinders is set to nm, a realistic value according to Ref. 24. The maximum length considered here is nm, corresponding to vertical rings in each cylindrical structure. The largest system considered in our simulations contains 132 840 Bchl c molecules, comparable with natural sizes. The positions and orientations of dipoles on each tubular wall are the same as in Complex A and B.
@@
Complex A - light-harvesting system subunit. The RC is directly surrounded by the so-called light harvesting complex I (LHI). LHI is in turn surrounded by several smaller light harvesting complexes, the LHII complexes 2, 55, 4. As shown in panel A of Fig. (2) two RCs are surrounded by the light-harvesting complex LHI B875, which is an S-shaped structure with 56 Bchl molecules. Then the LHI aggregate is surrounded by 10 LHII rings, each composed by 2 units, the B800 ring, composed by 9 Bchl molecules, and the B850 ring with 18 Bchl molecules. Both complexes LHI and LHII are J-aggregates. In particular LHI is a J-aggregate with two superradiant states close to the lowest excitonic state at 875 nm which are polarized in the ring plane. For the LHII aggregate, the B850 ring has two superradiant states at 850 nm, while the B800 ring has a main absorption peak at 800 nm. This hierarchical structure is able to absorb photons at different frequencies and to funnel the collected energy to the RC 56, 34, 4, 17, 16. In our simulation we model the RC as an aggregate of four BChls: two forming the tightly-coupled special pair and two accessory ones 17. Even if other molecules are present in the RC, these four Bchl molecules are the most relevant for the interaction with the electromagnetic field 57.
Complex B: the chromatophore. We model the whole 3D structure of the chromatophore using data from Ref. 34 collected by AFM, cryo-EM, X-ray crystallography and NMR measurements. In panel B of Fig. (2) our model of the chromatophore is shown with different substructures: i) LHII (blue rings); ii) LHI (orange structures); iii) RC, placed at the center of the LHI structures. Chromatophores, see Fig. (2 B), are connected to the cell membrane at their south pole, therefore the southern polar region is left empty (no molecular aggregates are placed there) to allow the contact with the cell membrane 58. The chromatophore spherical vesicles are reproduced by mapping small planar regions onto spherical ones containing the LHI-RCs or an LHII complex with the area-preserving inverse-Mollweide transformations, see Fig. (LABEL:PBprojection) in LABEL:AppA of the Supporting Information, used also in Ref. 34, 59, 60. While in nature the ratio between LHII and LHI-RC depends on illumination and other conditions, here we consider a vesicle model that contains S-shaped LHI-RCs and LHII complexes. The overall Bchl/RC ratio used in our simulation () is within the natural range for these vesicles, which is between and and corresponds to low-light growth conditions 34 where these systems show the largest efficiency. The data for the LHI, LHII and RC are taken from Ref. 17, 4. Even if it would be more realistic to add random rotations of LHII rings around the axis joining the center of the vesicle and the center of the LHII rings, we checked that the effect of such random rotations is small, see Fig. (LABEL:PBrotationLHII) of LABEL:AppE, and thus we did not include such rotations in our simulations.
@@ 2.3 The radiative Hamiltonian and the dipole approximation
Since photosynthetic antennae operate under natural sunlight, which is very dilute, the single-excitation approximation can be used, so that only states containing a single excitation have been considered. Choosing the basis states in the single-excitation manifold, where represents a state in which the molecule is excited while all the others are in the ground state, the systems can be described through a radiative non-Hermitian Hamiltonian which takes into account the interaction between the molecules mediated by the electromagnetic field (EMF) 13, 61, 12, 62. The radiative Hamiltonian reads:
@@ 2.4 Robustness to thermal noise and static disorder
Modelling the effect of noise and disorder in natural systems has been an important issue in the field of quantum biology. The environmental noise in the literature has been described with an open quantum system approach using an environment-system interaction even for non-Markovian baths 63, 64. Nevertheless due to the huge sizes considered in this manuscript we will adopt a minimal but solid modelling of noise and disorder. Noise due to thermal fluctuations will be taken into account by considering a state at thermal equilibrium, while disorder will be modelled as space dependent and time independent fluctuations of the site energies, keeping the couplings between the molecules constant. Both approaches have been widely used in literature 65, 28, 13, 8. They have the advantage to capture the main detrimental effects of noise and disorder even if a realistic modelling of the environment would require a much more sophisticated approach 53, 66, 67, 54, 27.
Given a quantum state specified by the density matrix it is possible to define its coherence length in the single-excitation manifold defined by the basis states 68, 69, 29 in the following way:
@@
is also able to produce a large thermal coherence length at room temperature. Note that even if we consider the coherence length at thermal equilibrium, this does not mean that out-of-equilibrium processes are not important in molecular nanotubes. Indeed in Ref. 70 strong evidence of ultra-fast transport in natural structures with transfer times less than 100 fs have been discussed. Nevertheless thermal equilibrium can be considered as a worst case scenario for coherences. For this reason assuming thermal equilibrium can be considered a good starting point to assess the structural robustness of quantum coherence to thermal noise.
@@
Natural complexes are not only affected by thermal noise but also by other sources of disorder due to the fluctuations in the local environment. In order to analyze the robustness to this kind of disorder, we have considered time-independent and space-dependent fluctuations of the excitation energy of the molecules in the aggregates. Specifically, we consider energy fluctuations which are uniformly distributed around the excitation energy of the molecules , between and , where represents the disorder strength. It is known that static disorder induces eigenmodes localization 71 and quenching of superradiance 65.
@@ 3 Numerical results
In this section we show the numerical results obtained for the antenna complexes of GSB and PB. Note that in LABEL:Spectra of Supporting Information the absorption spectra derived using our model are compared with the experimental results for GBS and PB, showing a good agreement.
@@ 3.1 Green Sulfur Bacteria complexes
Here, the complexes A, B, C of GSB described in Sec. 2.1 have been analyzed. In Fig. (3) the spectrum of the three complexes is shown for the maximal length considered, nm, corresponding to cylindrical aggregates with 180 rings. Thus, for the three complexes we have (complex A), (complex B), (complex C) chlorophyll molecules. Panels (A-C) in Fig. (3) show that the structure of all complexes allow for the emergence of red-shifted superradiant states, close to the lowest excitonic states. Moreover the amount of maximal superradiance () increases with the system size, see Fig. (4). The largest amount of superradiance is present in complex C which is the largest one and the closest to the natural size and structure of GSB antennae. Note that the fact that in larger aggregates superradiance is enhanced is far from being trivial. Indeed, if we would increase the system size by adding the molecules in the same positions but with randomized dipole directions, no superradiant enhancement would be present, as shown in LABEL:AppF-GSB of the Supporting Information.
@@
The differences between the NHH and the other Hamiltonian models is not due to the coupling strengths, which are almost identical in the three Hamiltonian models, as shown in Tab. (LABEL:tab-coupling) in LABEL:average-coupling, but it can be explained by the overlapping resonance criterion 36, see panel D in Fig. (3). Indeed, when the decay width becomes comparable with the eigenmodes mean level spacing, the non-Hermitian part of the Hamiltonian cannot be treated perturbatively and the dipole strength Eq. (6) does not describe anymore the decay widths of the eigenmodes.
@@
We now consider in Fig. (6), the three different complexes for a fixed length corresponding to 100 rings for each cylindrical aggregates ( nm), smaller than the maximal system length considered in Fig. (3,4). Also in this case we used the HH model to compute both figures of merit since for this length scale the HH model is quite accurate for all complexes, see discussion in LABEL:AppC of the Supporting Information. While the maximal dipole strength decreases with disorder Fig. (6 A), it is still much larger than one even for . On the other hand, the thermal coherence length shows a large robustness to disorder, Fig. (6 B) up to . In Fig. (6 C,D) the maximal dipole strength (panel C) and the thermal coherence length (panel D) are shown at and as a function of the number of chlorophyll molecules present in the three different complexes A,B,C. The most important feature of both figures of merit is that they grow as a larger portion of the photosynthetic antenna is considered, showing that the structure of the GSB photosynthetic antenna as a whole is able to support excitonic coherences. Note that in presence of disorder, the maximal dipole strength grows slower with compared with the case of zero disorder (panel C). On the hand, the value of the coherence length is the same for the two values of disorder strength considered, showing its robustness to disorder (panel D), but as increases it shows a tendency to saturate.
@@ 5 Conclusions
In analyzing different antenna complexes, we compared three distinct Hamiltonian models: the Dipole Hamiltonian, applicable in the small volume limit; the Hermitian Hamiltonian, valid when the decay widths are significantly smaller than the mean level spacing; and the full radiative non-Hermitian Hamiltonian model. For Purple bacteria chromatophores, all models yield consistent results due to the system’s small size (e.g., 60 nm) relative to the absorbed wavelength () and the limited number of Bchl molecules present. However, for the largest complexes considered in Green sulfur bacteria, only the non-Hermitian Hamiltonian model proved effective in providing reliable results about the eigenstates decay widths. Indeed consistent differences have been found in the largest complex analyzed (GSB complex C). The origin of such differences are not trivial. Indeed they are not due to the fact that the three Hamiltonian models considered (DH, HH, NHH) are characterized by different coupling strengths between the chlorophyll molecules, see LABEL:average-coupling, but to the fact that the decay widths in the largest system overlap leading to a different distribution of the dipole strengths among the eigenstates. This idea is supported by our analysis of the resonances overlap presented in this manuscript. This shows that the full non-Hermitian Hamiltonian is essential to analyze light matter interaction in molecular aggregates even in the small volume limit when resonances overlap.
The presence of very bright (superradiant) states in the whole photosynthetic antenna could shed new light in understanding the functionality of such complexes. In particular, the presence of superradiance and subradiance plays an important role for the efficiency of the energy transfer in these aggregates 72, 17. Let us also remark that the system dynamics strongly depends on the choice of the initial state which is mainly determined by the most superradiant states, see LABEL:AppD for a more detailed discussion about the delocalization of superradiant states in PB and GSB antennae.
@@
The cooperative properties, akin to those observed in photosynthetic aggregates, are inspiring numerous proposals for engineering artificial devices for light-harvesting and clean energy production 73, 74, 75, 76, 77, 78, 79, 80. Thus, our study could significantly impact the development of quantum devices for photon detection and light harvesting.
@@
Additional details and methods, including further numerical and experimental data are available in the Supporting Information.